OpenStudy (anonymous):

solve quadratic equation x^2-10x-24=0

OpenStudy (anonymous):

gotta factor

OpenStudy (anonymous):

show me your steps to see if i did it right?

OpenStudy (anonymous):

but do you know how to factor?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have (x-12)(x+2)=0, then x-12=0 and x+2 =0 x=12 and x=-2

OpenStudy (anonymous):

x=-2,12

OpenStudy (anonymous):

test both values out in the original equation

OpenStudy (anonymous):

if they both equal 0 then this answer is correct

OpenStudy (anonymous):

thats what i got i just wasnt positive but thanks very much

OpenStudy (anonymous):

np